OpenStudy (anonymous):

3/2x=7/5 solve the proportion.?

OpenStudy (anonymous):

x=15/14 just cross multiply and solve for x

OpenStudy (anonymous):

\[3\div 2x = 7\div 5\] \[x =7\times 2\div 5 \times 3\] \[x = 14 \div 15\] \[x = 0.93\]

OpenStudy (anonymous):

oooh okay i get it now thanx

OpenStudy (anonymous):

then please award a medal noe